OpenStudy (anonymous):

Solve: 2(x + 2) − (x + 1) = 8 2 3 4 5

OpenStudy (anonymous):

We have : 2*(x)+2*(2)=2x+4 and have : 2x+4-x-1=x-3 k ?!

OpenStudy (anonymous):

and have this equation : x-3=8

OpenStudy (anonymous):

So it would be 5

OpenStudy (anonymous):

@E.ali should it be x+3 right?

OpenStudy (anonymous):

since 2x+4 -x-1 = x+3?

OpenStudy (anonymous):

Yep ! And have : x+3=8 x=8-3

OpenStudy (anonymous):

1x + 3 = 8 subtract the 3 from 8. your answer is 5.